# Asymptotic independence in higher dimensions and its implications on risk management

### v2

@@ 4 $k$-wise asymptotic independence
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with survival copula $\widehat{C}$ is $k$-wise asymptotically independent for a fixed $k\in\{2,\ldots,d\}$, if for all $S\subseteq\mathbb{I}_{d}$ with $2\leq|S|\leq k$, we have
@@
Note that for a $d$-dimensional copula, $d$-wise asymptotic independence is the same as mutual asymptotic independence (and of course 2-wise is the same as pairwise). Again, following Proposition 3.2, we may check that mutual asymptotic independence indeed implies $k$-wise asymptotic independence for all $k\in\{2,\ldots,d\}$. The reverse of the previous implication is, of course, not true; the examples in the following section also show this.
@@ 4.1 Examples
Example 4.3 (ACIG copula).
This Archimedean copula based on the Laplace transform (LT) of an Inverse Gamma distribution, called ACIG copula in short, was introduced in Hua and Joe [23], i.e, if $Y=X^{-1}$ and $X\sim\text{Gamma}(\alpha,1)$ for $\alpha>0$, then the generator of this Archimedean copula is given by the LT of $Y$. The expression of the generator includes the Bessel function of the second kind. Closed-form expressions of the copula $C^{\phi}$ and survival copula $\widehat{C}^{\phi}$ are not easy to write down; nevertheless, from computations in Hua and Joe [23, Example 4] and Hua, Joe and Li [24, Example 4.4], we can conclude that for any $d\geq 2$, the survival copula of the ACIG copula with parameter $\alpha>0$ has the following asymptotic behavior:
@@
    If $0<\alpha\leq 1$, then it exhibits asymptotic upper tail dependence.
@@
    If $11<\alpha\leq d-$ 1, then it exhibits pairwise asymptotic independence but not mutual asymptotic independence. If additionally $k-1<\alpha\leq k$ for $k\in\{2,\ldots,d-1\}$, then the ACIG still exhibits $i$-wise asymptotic independence for all $i\in\{2,\ldots,k\}$, but not $(k+1)$-wise asymptotic independence.
@@
    If $\alpha>d-1$, then it exhibits $k$-wise asymptotic independence for all $k\in\{2,\ldots,d\}$, and hence mutual asymptotic independence as well.
Example 4.4 (Gaussian copula).
The Gaussian copula has been popular in modeling dependence in a wide variety of applications. It turns out that a class of Gaussian copula models is also able to capture the presence of $k$-wise asymptotic independence and not $(k+1)$-wise asymptotic independence. This is demonstrated in the following result.
@@
For any $k\in\{2,\ldots,d-1\}$ and $S_{1},S_{2}\subseteq\{1,\ldots,d\}$ with $|S_{1}|=k$ and $|S_{2}|=k+1$, there exists a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with a Gaussian copula $C^{\Sigma}$ and a positive-definite correlation matrix $\Sigma$, such that $\boldsymbol{Z}$ exhibits $k$-wise asymptotic independence but not $(k+1)$-wise asymptotic independence and for any $x>0$,
@@
This theorem not only provides the existence of a $k$-wise asymptotically independent Gaussian copula, but it also gives the striking feature of the copula behavior in (4.2) where, surprisingly the value of $x$ has no influence. For the derivation of worst-case measures for risk contagion under distributional ambiguity in the next section, this result turns out to be indispensable.
@@
It is important to mention here that although most popular copula families are bivariate in nature, Joe [25, Chapter 4] lists multiple extensions of bivariate copulas to general high dimensions; many such copulas can be explored for creating models with particular levels of asymptotic independence as necessitated by the context.
@@ 5 Implication on risk management under distributional ambiguity
In financial risk management, a variety of risk measures are used to assess the risk contagion between different financial products including stocks, bonds and equities. Such contagion or systemic risk measures are often based on conditional probabilities and range from computing regular conditional tail probabilities to CoVaR, marginal expected shortfall (MES), marginal mean excess (MME), and more; see[8, 1] for details.
@@
the $(1-\gamma)$-quantile of $Z$ where $\inf\emptyset:=\infty$ (cf. [16]). If $Z\sim F$ is continuous, then $F$ is invertible with inverse $F^{-1}$ and for $10<\gamma<$ 1 we have $\mathrm{VaR}_{\gamma}(Z)=F^{-1}(1-\gamma)$.
Consider the returns from a portfolio of $d>1$ stocks being given by the random vector $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})\sim F$. Suppose we are interested in measuring the risk of $Z_{1}$ having an extremely large value, given that all variables in some non-empty subset $J\subset\mathbb{I}_{d}\backslash\{1\}$ with $|J|=\ell$ are at extremely high levels. This can be captured via the following conditional tail probability
@@
The risk measure CoVaR was introduced in the bivariate setting for $J={2}$ to capture risk contagion, as well as systemic risk by Adrian and Brunnermeier [1] where they used the conditioning event to be $Z_{2}=\mathrm{VaR}_{\gamma_{2}}(Z_{2})$; this was later modified by Girardi and Ergün [17] to $Z_{2}>\mathrm{VaR}_{\gamma_{2}}(Z_{2})$ with the restriction that $\gamma_{1}=\gamma_{2}$; this latter definition has been widely used in dependence modeling ([35, 32, 19, 12]) and is generalized in our definition given in (5.3).
In risk management applications, computing quantities like CTP and CoVaR requires knowledge of the joint distribution of the risk vector $\boldsymbol{Z}$. Even if the univariate distributions of all the marginal variables can be estimated, the joint distribution often remains unknown and relatively more involved for estimation purposes. An approach often used is to provide a worst-case value for such risk measures under certain constraints on the joint distribution of the variables. Naturally, for such tail risk measures, constraints can be provided in terms of their joint asymptotic tail behavior, including pairwise, mutual or $k$-wise asymptotic independence. It turns out that under different constraints, we may obtain different tail behavior for the worst-case measures. To further this discussion, let us define $\mathcal{P}$ to be the class of all probability distributions in $\mathbb{R}^{d}$ with continuous marginal distributions.
Definition 5.1.
For $k\in\{2,\ldots,d\}$ define the classes of distributions
@@
and similarly the restrictions to distributions with Gaussian copulas
$\mathcal{N}_{k}$ $:=\mathcal{P}_{k}\cap\{F\in\mathcal{P}:F\text{ has a Gaussian copula }C^{\Sigma}\text{ with }\Sigma\text{ positive definite}\}.$
@@
Since the joint distributions are unknown we may want to find the worst case $\mathrm{CTP}$ or CoVaR in such cases where $F\in\mathcal{P}_{k}\subset\mathcal{P}$ or $F\in\mathcal{N}_{k}$, $k\in\{2,\ldots,d\}$. First, we present the result for the $\mathrm{CTP}$.
@@
Let $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})\sim F$ where $d\geq 2$ and all the variables have continuous marginal distributions. Furthermore, suppose $J\subset\mathbb{I}_{d}\backslash\{1\}$ with $|J|=\ell$.
@@
The results indicate a qualitatively different behavior of the worst-case CTP depending on whether the tail dependence exhibits $k$-wise asymptotic independence with $k>|J|$ vis-a-vis $k\leq|J|$. When $k>|J|$, $\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})$ converges to 00 as $\gamma\downarrow 0$, suggesting that extreme large losses of $Z_{j}$ for all $j\in J$ have a negligible influence on extreme large losses of $Z_{1}$. In contrast, when $k\leq|J|$, there exists a $k$-wise asymptotically independent distribution function $F$, which is as well pairwise asymptotically independent, such that extreme large losses of $Z_{j}$ for all $j\in J$ result, with a probability converging to $1$, in an extremely large loss of $Z_{1}$. In particular, for the Gaussian copula that is an astonishing result because it is in contrast to the belief that there are no joint extremes. This shows that for measuring risk contagion it is important to distinguish between these different concepts of tail independence and assuming an improper notion of asymptotic independence for our risk portfolio may lead to either underestimation or overestimation of the risk contagion.
In the following, we investigate the asymptotic behavior of the measure $\mathrm{CoVaR}$. For technical reasons, we restrict the class $\mathcal{P}_{k}$ slightly, in particular we will assume that $F_{1}$, the distribution of $Z_{1}$, is Pareto distributed, i.e., $F_{1}(z)=1-z^{-\alpha}$, $z\geq 1$, for some $\alpha>0$.
Definition 5.3.
Suppose $\mathcal{P}^{*}:=\{F\in\mathcal{P}:F_{1}\text{ is Pareto distributed}\}$. For $k\in\{2,\ldots,d\}$ define the classes
@@
Instead of assuming that $F_{1}$ follows a Pareto distribution, it is possible to consider a broader class, allowing $F_{1}$ to have a regularly varying tail. However, this approach makes the proofs more technical without providing any further valuable insights, hence we have exhibited our results for the smaller class $\mathcal{P}_{k}^{*}$ for the purpose of exposition.
@@
Let $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})\sim F$ where $d\geq 2$ and all the variables have continuous marginal distributions and $F_{1}$ is a Pareto distribution. Furthermore, let $J\subset\mathbb{I}_{d}\backslash\{1\}$ with $|J|=\ell$.
@@
Akin to the case of finding for the worst-case CTP, we observe that the worst-case CoVaR also has a qualitatively different behavior depending if the tail dependence exhibits $k$-wise asymptotic independence with $k>|J|$, or with $k\leq|J|$. When $k>|J|$, the ratio $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})/{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ converges to 00, reflecting that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ increases at a negligible rate in comparison to $\mathrm{VaR}_{\gamma_{2}}(Z_{1})$ as $\gamma_{2}\downarrow 0$ and that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ is relatively small, i.e, the required risk reserve capital is low. But if $k\leq|J|$, there exists a $F\in\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$ where $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})/{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ converges to $\infty$, so that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ may increase much faster to $\infty$ than $\mathrm{VaR}_{\gamma_{2}}(Z_{1})$ as $\gamma_{2}\downarrow 0$, giving a relatively high $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ and a higher reserve risk capital requirement.
@@ 6 Conclusion
In this paper, we provide a notion of truly multivariate asymptotic independence that is useful in comparing extreme events in different dimensions beyond mere pairwise comparisons. This parallels the dichotomy of mutual independence vis-a-vis pairwise independence for multivariate random vectors. We believe this new notion also provides an alternate pathway for characterizing extremal dependence for high-dimensional problems relating to tail events. We have illustrated using examples of particular copula models, including a few from the Archimedean family along with the Gaussian and the Marshall-Olkin copula. The copulas considered often exhibit at least pairwise asymptotic independence if not mutual asymptotic independence. For both Archimedean and Gaussian copulas, we presented examples exhibiting not only mutual asymptotic independence but as well exhibiting only pairwise asymptotic independence but not mutual asymptotic independence. In particular, for the Gaussian copula, this result is quite striking since it is in contrast to the common belief that the Gaussian copula does not allow joint extremes. We have also introduced the concept of $k$-wise asymptotic independence which generalizes the two notions and brings them under the same umbrella. Here we have shown that for any $k\in\{2,\ldots,d\}$ there exists a $k$-wise asymptotic independent Gaussian copula (which is not $(k+1)$-wise asymptotically independent if $k<d$). Assumptions of different notions of asymptotic tail independence significantly impact measures of risk contagion within a financial system, such as conditional tail probabilities (CTP) or contagion value-at-risk (CoVaR), depending on the specific context. Overlooking these concepts and assuming only pairwise asymptotic independence for models could lead to an underestimation of risks.
Appendix A Proofs
A.1 Proofs of Section 2
Proof of Proposition 2.6.
Consider an MO survival copula with parameter set $\Lambda=\{\lambda_{S}:\emptyset\neq S\subseteq\mathbb{I}_{d}\}$. We need to check (2.2) for any distinct $j,\ell\in\mathbb{I}_{d}$. W.l.o.g. assume $j=1,\ell=2$. Then
$\widehat{C}^{\mathrm{MO}(\Lambda)}_{\{1,2\}}(u,u)$ $=\widehat{C}^{\mathrm{MO}(\Lambda)}(u,u,1,\ldots,1)=u^{\eta^{*}}$
where
$\eta^{*}$ $=\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S,2\notin S}\eta_{1}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\eta_{2}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1,2\in S}\max\{\eta_{1}^{S},\eta_{2}^{S}\}$
$=\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S,2\notin S}\frac{\lambda_{S}}{\Delta_{1}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\frac{\lambda_{S}}{\Delta_{2}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1,2\in S}\max\left\{\frac{\lambda_{S}}{\Delta_{1}},\frac{\lambda_{S}}{\Delta_{2}}\right\}$
with $\Delta_{j}=\sum\limits_{J\subseteq\mathbb{I}_{d}\atop j\in J}\lambda_{S}$ for $j=1,2$. W.l.o.g. assume $\Delta_{1}\geq\Delta_{2}$. Then
$\eta^{*}$ $\geq\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S,2\notin S}\frac{\lambda_{S}}{\Delta_{1}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\frac{\lambda_{S}}{\Delta_{1}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1,2\in S}\frac{\lambda_{S}}{\Delta_{1}}$
$=\frac{1}{\Delta_{1}}\Big{(}\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S}\lambda_{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\lambda_{S}\Big{)}$
$=\frac{1}{\Delta_{1}}\Big{(}\Delta_{1}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\lambda_{S}\Big{)}>1$
since $\lambda_{S}>0$ for all non-empty $S$. Hence,
$\widehat{C}^{\mathrm{MO}(\Lambda)}_{1,2}(u,u)$ $=u^{\eta^{*}}=o(u),\quad\quad u\downarrow 0.$
showing (2.2). ∎
A.2 Proofs of Section 3
Proof of Proposition 3.6.
Let $S\subseteq\mathbb{I}_{d}$ such that $|S|=k\geq 2$. W.l.o.g. let $S=\{1,\ldots,k\}$. Fix any $\ell\in S$. For notational convenience, we will take w.l.o.g. $\ell=k$. Then for $0\leq u\leq 1$,
$\frac{\widehat{C}_{S}^{\mathrm{MO}(\Lambda)}(u,\ldots,u)}{\widehat{C}_{S\setminus\{k\}}^{\mathrm{MO}(\Lambda)}(u,\ldots,u)}$ $=\frac{\widehat{C}^{\mathrm{MO}(\Lambda)}(\overbrace{u,\ldots,u}^{k},\overbrace{1,\ldots,1}^{(d-k)})}{\widehat{C}^{\mathrm{MO}(\Lambda)}(\underbrace{u,\ldots,u}_{(k-1)},\underbrace{1,\ldots,1}_{(d-k+1)})}$
$=\prod_{i=1}^{d}\prod_{\begin{subarray}{c}J\subset\mathbb{I}_{d}:\\ |J|=i,k\in J\end{subarray}}\bigwedge_{j\in J,j\leq k}u^{\eta_{j}^{J}}$
$=u^{\eta_{k}^{\{k\}}}\Bigg{(}\prod_{i=2}^{d}\prod_{\begin{subarray}{c}J\subset\mathbb{I}_{d}:\\ |J|=i,k\in J\end{subarray}}\bigwedge_{j\in J,j\leq k}u^{\eta_{j}^{J}}\Bigg{)}$
[…]
The second equality above is obtained after canceling all common terms appearing in both the numerator and denominator; these canceled terms correspond to the inner/second product not containing the $k$-th co-ordinate in the set $J$. Hence, all terms in the denominator are canceled. The third equality is obtained by accounting for $i=1$ in the outer product separately. By model definition $\eta_{k}^{\{k\}}>0$ and $g(u)=u^{\beta}$ for some $\beta\geq 0$. Hence
$\lim_{u\downarrow 0}\frac{\widehat{C}_{S}^{\mathrm{MO}(\Lambda)}(u,\ldots,u)}{\widehat{C}_{S\setminus\{k\}}^{\mathrm{MO}(\Lambda)}(u,\ldots,u)}=\lim_{u\downarrow 0}u^{\eta_{k}^{\{k\}}+\beta}=0,$
showing (3.1) and hence proving the claim. ∎

### v3

@@ 4 $k$-wise asymptotic independence
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions and survival copula $\widehat{C}$ is $k$-wise asymptotically independent for a fixed $k\in\{2,\ldots,d\}$, if for all $S\subseteq\mathbb{I}_{d}$ with $2\leq|S|\leq k$, we have
@@
Remark 4.2.
The concept of $k$-wise asymptotic independence is a measure of dependence in the extremes. If a random vector $\boldsymbol{Z}$ exhibits $k$-wise asymptotic independence but not $(k+1)$-wise asymptotic independence, then there exists a combination of at exactly $k$ components in $\boldsymbol{Z}$, so that when these are large another additional component is large as well. Fewer than $k$ large components cannot produce a large value in another component. Consequently, lower values of $k$ reflect a stronger dependence in the extremes.
Note that for any $d$-dimensional copula, $d$-wise asymptotic independence is the same as mutual asymptotic independence (and of course 2-wise is the same as pairwise). Again, following Proposition 3.3, we may check that mutual asymptotic independence indeed implies $k$-wise asymptotic independence for all $k\in\{2,\ldots,d\}$. The converse of the previous implication is, of course, not true; the examples in the following section also show this.
@@ 4.1 Examples
4.1.1 ACIG copula
This Archimedean copula based on the Laplace transform (LT) of an Inverse Gamma distribution, called ACIG copula in short, was introduced in Hua and Joe [24], i.e, if $Y=X^{-1}$ and $X\sim\text{Gamma}(\alpha,1)$ for $\alpha>0$, then the generator of this Archimedean copula is given by the LT of $Y$. The expression of the generator includes the Bessel function of the second kind. Closed-form expressions of the copula $C^{\phi}$ and survival copula $\widehat{C}^{\phi}$ are not easy to write down; nevertheless, from computations in Hua and Joe [24, Example 4] and Hua, Joe and Li [25, Example 4.4], we can conclude that for any $d\geq 2$, the survival copula of the ACIG copula with parameter $\alpha>0$ has the following asymptotic behavior:
@@ 4.1.1 ACIG copula
    If $0<\alpha\leq 1$, the ACIG copula exhibits asymptotic upper tail dependence.
@@
    If $1<\alpha\leq d-1$, the ACIG copula exhibits pairwise asymptotic independence but not mutual asymptotic independence. If additionally $k-1<\alpha\leq k$ for $k\in\{2,\ldots,d-1\}$, then the ACIG copula still exhibits $i$-wise asymptotic independence for all $i\in\{2,\ldots,k\}$, but not $(k+1)$-wise asymptotic independence.
@@
    If $\alpha>d-1$, the ACIG copula exhibits $k$-wise asymptotic independence for all $k\in\{2,\ldots,d\}$, and hence mutual asymptotic independence as well.
4.1.2 Gaussian copula
The Gaussian copula has been popular in modeling dependence in a wide variety of applications. It turns out that a class of Gaussian copula models is also able to capture the presence of $k$-wise asymptotic independence and not $(k+1)$-wise asymptotic independence. This is demonstrated in the following result, whose proof is given in Appendix B.
@@ 4.1.2 Gaussian copula
Suppose $k\in\{2,\ldots,d-1\}$ and $S_{1}\subseteq S_{2}\subseteq\{1,\ldots,d\}$ with $|S_{1}|=k$ and $|S_{2}|=k+1$. Then there exists a Gaussian copula $C^{\Sigma}$ and a positive-definite correlation matrix $\Sigma$, such that $C^{\Sigma}$ exhibits $k$-wise asymptotic independence but not $(k+1)$-wise asymptotic independence and for any $x>0$,
@@
This theorem not only provides the existence of a $k$-wise asymptotically independent Gaussian copula, but it also gives the striking feature of the copula behavior in (4.2) where, surprisingly the value of $x$ has no influence. This means that for a random vector $\boldsymbol{Z}$ with Gaussian copula $C^{\Sigma}$, as given in Theorem 4.4, and identically distributed marginals, large values in all the $S_{1}$ components result in an extremely large value in the single component of $S_{2}\backslash S_{1}$; hence there is a strong dependence between the extremes of the components of $S_{1}$ and that of $S_{2}\backslash S_{1}$. All components in $S_{1}$ must be large at the same time; only a few components that are large do not result in a large value in the $S_{2}\backslash S_{1}$ component. This is demonstrated quite nicely in the following example, which was used in the proof of Theorem 4.4.
Example 4.5.
Let the correlation matrix of the Gaussian copula be given as
$\Sigma=\begin{bmatrix}\boldsymbol{I}_{d-1}\;\;\;&\rho\boldsymbol{1}_{d-1}\\[10.00002pt] \rho\boldsymbol{1}_{d-1}^{\top}&1\end{bmatrix}$
for some $\rho\in(0,1)$. Then for $\rho\in(1/(k-1),1/(k-2))$ the Gaussian copula $\widehat{C}^{\Sigma}$ is (k1)(k-1)-wise asymptotically independent but not $k$-wise asymptotically independent (see proof of Theorem 4.4); therefore, if the first (k1)(k-1) components are jointly large then as well the last component is large. But if we consider fewer components than (k1)(k-1) components to be large, they have no effect on the size of the last component. Note that here $k-1=\lceil\rho^{-1}\rceil=\inf\{m\in\mathbb{N}:\rho^{-1}\leq m\}$. Thus, for a high value of $\rho$, fewer components, namely only the first $\lceil\rho^{-1}\rceil$ components, result in a large value in the last component. Although $\rho$ provides a measure of linear dependence, in this example, it also measures the dependence in the extremes which is reflected in the degree $\lceil\rho^{-1}\rceil$ of asymptotic independence.
For the derivation of worst-case measures for risk contagion under distributional ambiguity in the next section, Theorem 4.4 turns out to be indispensable.
@@
It is important to mention here that although most popular copula families are bivariate in nature, Joe [26, Chapter 4] lists multiple extensions of bivariate copulas to general high dimensions; many such copulas can be explored for creating models with particular levels of asymptotic independence as necessitated by the context.
@@ 5 Implication on risk management under distributional ambiguity
In financial risk management, a variety of risk measures are used to assess the risk contagion between different financial products, including stocks, bonds and equities. Such contagion or systemic risk measures are often based on conditional probabilities and range from computing regular conditional tail probabilities to CoVaR, marginal expected shortfall (MES), marginal mean excess (MME), and more; see [8, 1] for details.
@@
the $(1-\gamma)$-quantile of $Z$ where $\inf\emptyset:=\infty$ (cf. [16]). If $Z\sim F$ is continuous, then $F$ is invertible with inverse $F^{-1}$ and for $0<\gamma<1$ we have $\mathrm{VaR}_{\gamma}(Z)=F^{-1}(1-\gamma)$.
[…]
@@
The risk measure CoVaR was introduced in the bivariate setting for J=2J={2} to capture risk contagion, as well as systemic risk by Adrian and Brunnermeier [1] where they used the conditioning event to be $Z_{2}=\mathrm{VaR}_{\gamma_{2}}(Z_{2})$; this was later modified by Girardi and Ergün [18] to $Z_{2}>\mathrm{VaR}_{\gamma_{2}}(Z_{2})$ with the restriction that $\gamma_{1}=\gamma_{2}$; this latter definition has been widely used in dependence modeling [37, 33, 20, 12] and is generalized in our definition given in (5.3).
In risk management applications, computing quantities like CTP and CoVaR requires knowledge of the joint distribution of the risk vector $\boldsymbol{Z}$. Even if the univariate distributions of all the marginal variables can be estimated, the joint distribution often remains unknown and relatively more involved for estimation purposes. An approach often used is to provide a worst-case value for such risk measures under certain constraints on the joint distribution of the variables. Naturally, for such tail risk measures, constraints can be provided in terms of their joint asymptotic tail behavior, including pairwise, mutual or $k$-wise asymptotic independence. It turns out that under different constraints, we may obtain a different tail behavior for the worst-case measures. To further this discussion, let us define $\mathcal{P}$ to be the class of all probability distributions in $\mathbb{R}^{d}$ with continuous marginal distributions. For $k\in\{2,\ldots,d\}$ define the classes of distributions
$\mathcal{P}_{k}$ $:=\{F\in\mathcal{P}:F\text{ possesses $k$-wise asymptotic independence}\},$
and similarly, the restrictions to distributions with Gaussian copulas
@@
Since the joint distributions are unknown we may want to find the worst case $\mathrm{CTP}$ or CoVaR in such cases where $F\in\mathcal{P}_{k}\subset\mathcal{P}$ or $F\in\mathcal{N}_{k}$, $k\in\{2,\ldots,d\}$. First, we present the result for the $\mathrm{CTP}$. The proof of this theorem and all subsequent results in this section are given in Appendix C.
@@
[…]
@@
The results indicate a qualitatively different behavior of the worst-case CTP depending on whether the tail dependence exhibits $k$-wise asymptotic independence with $k>|J|$ vis-a-vis $k\leq|J|$. When $k>|J|$, $\mathrm{CTP}_{\gamma}(\boldsymbol{Z}_{1|J})$ converges to 0 as $\gamma\downarrow 0$, suggesting that extreme large losses of $Z_{j}$ for all $j\in J$ have a negligible influence on extreme large losses of $Z_{1}$. In contrast, when $k\leq|J|$, there exists a $k$-wise asymptotically independent distribution function $F$, which is also pairwise asymptotically independent, such that extremely large losses of $Z_{j}$ for all $j\in J$ result, with a probability converging to $1$, in an extremely large loss of $Z_{1}$. In particular, for the Gaussian copula that is an astonishing result because it is in contrast to the belief that there are no joint extremes. This shows that for measuring risk contagion it is important to distinguish between these different concepts of tail independence and assuming an improper notion of asymptotic independence for our risk portfolio may lead to either underestimation or overestimation of the risk contagion.
In the following, we investigate the asymptotic behavior of the measure $\mathrm{CoVaR}$. For technical reasons, we restrict the class $\mathcal{P}_{k}$ slightly, in particular we will assume that $F_{1}$, the distribution of $Z_{1}$, is Pareto distributed, i.e., $F_{1}(z)=1-z^{-\alpha}$, $z\geq 1$, for some $\alpha>0$. Suppose $\mathcal{P}^{*}:=\{F\in\mathcal{P}:F_{1}\text{ is Pareto distributed}\}$. For $k\in\{2,\ldots,d\}$ define the classes
@@
Instead of assuming that $F_{1}$ follows a Pareto distribution, it is possible to consider a broader class, allowing $F_{1}$ to have a regularly varying tail. However, this approach makes the proofs more technical without providing any further valuable insights, hence, we have exhibited our results for the smaller class $\mathcal{P}_{k}^{*}$ for the purpose of exposition.
@@
[…]
@@
Akin to the case of finding for the worst-case CTP, we observe that the worst-case CoVaR also has a qualitatively different behavior depending if the tail dependence exhibits $k$-wise asymptotic independence with $k>|J|$, or with $k\leq|J|$. When $k>|J|$, the ratio $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})/{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ converges to 0, reflecting that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ increases at a negligible rate in comparison to $\mathrm{VaR}_{\gamma_{2}}(Z_{1})$ as $\gamma_{2}\downarrow 0$ and that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ is relatively small, i.e, the required risk reserve capital is low. But if $k\leq|J|$, there exists a $F\in\mathcal{N}_{k}^{*}\subseteq\mathcal{P}_{k}^{*}$ where $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})/{\mathrm{VaR}_{\gamma_{2}}(Z_{1})}$ converges to $\infty$, so that $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ may increase much faster to $\infty$ than $\mathrm{VaR}_{\gamma_{2}}(Z_{1})$ as $\gamma_{2}\downarrow 0$, giving a relatively high $\text{CoVaR}_{\gamma_{1},\gamma_{2}}(\boldsymbol{Z}_{1|J})$ and a higher reserve risk capital requirement.
@@ 6 Conclusion
In this paper, we provide a notion of multivariate asymptotic independence that is useful in comparing extreme events in different dimensions beyond mere pairwise comparisons, which has been traditionally used in the literature. This parallels the dichotomy of mutual independence vis-a-vis pairwise independence for multivariate random vectors. We believe this new notion also provides an alternate pathway for characterizing extremal dependence for high-dimensional problems relating to tail events. We have illustrated using examples of particular copula models, including a few from the Archimedean family along with the Gaussian and the Marshall-Olkin copula. The copulas considered often exhibit at least pairwise asymptotic independence if not mutual asymptotic independence. For both Archimedean and Gaussian copulas, we presented examples exhibiting not only mutual asymptotic independence but also exhibiting only pairwise asymptotic independence but not mutual asymptotic independence. In particular, for the Gaussian copula, this result is quite striking since it is in contrast to the common belief that the Gaussian copula does not allow joint extremes. We have also introduced the concept of $k$-wise asymptotic independence which generalizes these two notions (pairwise and mutual) and brings them under the same umbrella. Here we have shown that for any $k\in\{2,\ldots,d\}$ there exists a $k$-wise asymptotically independent Gaussian copula (which is not $(k+1)$-wise asymptotically independent if $k<d$). Moreover, we have exhibited that these assumptions of different notions of asymptotic tail independence significantly impact measures of risk contagion within a financial system, such as conditional tail probabilities (CTP) or Contagion Value-at-Risk (CoVaR), depending on the specific context. Overlooking these concepts and assuming merely pairwise asymptotic independence for models may often lead to a significant underestimation of risks.
Appendix A Proofs of Section 3